OpenStudy (mikezack123):

Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 − 25 = 0 and x2 = 2x + 15 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (mikezack123):

i did part 1 and 2 i need 3

OpenStudy (mikezack123):

here's my work Part 1: 1st equation x^2−25=0 x^2=25 x=−5 or x=5 2nd equation second one x^2=2x+15 x^2−2x−15=0 x=5 or x=-3 Part 2: The zeros are where the graph of y=x2−25 cross the x axis same for y=x^2−2x−15 Part 3:

OpenStudy (mikezack123):

@satellite73 @jim_thompson5910

OpenStudy (mikezack123):

@ganeshie8 @myininaya @Luigi0210